OpenStudy (anonymous):

If A = 1 and B = 0, then what is the value of (NOT A) OR B?

OpenStudy (anonymous):

(NOT A) evaluates to => 0 0 or 0 evaluates to => ?

OpenStudy (anonymous):

0 + 0 = 0 1 + 0 = 1 0 + 1 = 1 1 + 1 = 0

OpenStudy (anonymous):

What does 0 or 0 evaluate to

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

You need to recall some definitions. 0 AND 0 => 0 1 AND 0 => 0 0 AND 1 => 0 1 AND 1 => 1 0 OR 0 => 0 0 OR 1 => 1 1 OR 0 => 1 1 OR 1 => 1 NOT(0) => 1 NOT(1) => 0

OpenStudy (anonymous):

Use the table to make associates ex: AND (if they are both true => true otherwise false) OR (if at least one is true => true otherwise false) NOT (if true => false otherwise true)